1) A current of 1.7A is passed through 300mm of 0.16mol/dm^3 of ZnSO4 for 230 secs with a current deficiency of 90%.Find the molarity of Zn2+ after the deposition of Zn. Assume the vol of solution remains constant during electrolysis.

2) A 300 ml solution of NACl was electrolysed for 6mins. If the pH of the final solution was12.24, calculate the average current used?







1
Expert's answer
2021-03-10T06:57:05-0500

1)"Q= it"

Q=1.7× 230=391c

1mole = 96500c

391/96500= 0.0041moles of zinc

Molarity= moles/ 1000ml

= 0.41× 10-5